# Invisible Walls in Cities: Leveraging Large Language Models to Predict Urban Segregation Experience with Social Media Content

### v3

@@ 6 User Study
Figure 3: Accuracy of visitor attraction prediction.
@@
Figure 4 presents the results of the summary preference evaluation. Across all sampled POIs, 60%-80% of participants prefer the codebook-guided summary over the vanilla one, with an average improvement of 40%, showing that our approach more effectively distills useful information to support human judgment. As Appendix D illustrates, the codebook-guided summaries capture nuanced insights on the potential deterrants for certain minority groups, which are often absent or less explicit without codebook guidance.
7 Prediction Experiments
7.1 Experiment Settings
We compare our model against four baselines. The first baseline is an MLP making use of only the population information. The other three baselines process the social media content with a frozen pre-trained text embedding model, and fuse it with population information. We adopt three widely-used powerful embedding models: GTE-base Li et al. (2023), BERT-base Devlin et al. (2019), and GloVE-330B Pennington et al. (2014). The structures of the population and embedding MLPs are identical to those in our model. We use four metrics to evaluate model performances: mean squared error (MSE), root mean squared error (RMSE), mean absolute error (MAE), and coefficient of determination (R2).
We employ GPT-4o for reflective coding on a small subset (n=190) of POIs, and GPT-4o mini for codebook-guided reasoning on all POIs for economic considerations. All models are trained using the Adam optimizer to minimize the MSE loss, with hyperparameters (e.g., learning rate, weight decay) tuned based on validation performances. POIs in each city are randomly split into training, validation, and test sets with a ratio of 6:2:2. For further implementation details, please refer to Appendix C and our released code  [^5] https://github.com/LinChen-65/REEM.
@@ 7.1 Experiment Settings
Table 3: Ablation study.
@@
Our Model 0.0087±0.0005 0.0929±0.0028 0.0745±0.0024 0.2781±0.0424
Table 4: Cross-city generalization performances.
Figure 4: Summary preference evaluation.
7.2 Experiment Results
Table 2 presents the performances of our RE’EM model alongside four baselines, with results reported as the mean and standard deviation over five repetitions on 50% randomly sampled test sets. A key insight is that all models incorporating social media data outperform the population-only baseline. This result underscores the predictive power of social media content, which provides crucial insights beyond static population distributions. It supports our hypothesis that segregation patterns are shaped not only by demographic distributions but also by the social and cultural perceptions reinforcing the “invisible walls” in cities. Further, RE’EM consistently and significantly surpasses all embedding-based baselines. It achieves a 9.33%, 4.75%, and 3.07% improvement in MSE, RMSE, and MAE, respectively, over the population-only baseline. R2 increases to 0.3885, marking a substantial 22.79% improvement. These results indicate that while social media data encapsulates rich and valuable information, a naive embedding strategy is insufficient to fully capture the interplay between population distribution, POI attributes, and visitor dynamics. In contrast, the structured multi-view approach of RE’EM effectively integrates heterogeneous information sources, enabling accurate predictions of segregation.
To validate our design of different model components, we perform an ablation study. As Table 3 shows, each channel (reasoning, embedding, population) exhibits certain predictive capabilities, but none achieves optimal performance. The full RE’EM model, which strategically combines all three channels, consistently delivers superior results. Thus, the integration of multi-view signals enables complementary information flow, allowing the model to refine its predictions beyond what any single feature source can achieve.
To assess the generalization capability across cities, we use the codebook obtained in Philadelphia to train models in the other three cities. As shown in Table 4, RE’EM consistently achieves the best performance in all three cities along all metrics. Compared to the strongest baseline, RE’EM achieves substantial improvements, with MSE reduced by as much as 17.02%, MAE reduced by as much as 9.64%, and R² increasing by as much as 63.71%. These findings indicate that RE’EM, and specifically, our constructed codebook, captures fundamental segregation mechanisms that extend beyond city-specific characteristics, making it a promising tool across diverse geographic and socio-economic settings.
8 Discussion
@@ Appendix A Variability of POI features
Figure 5: The Coefficient of Variation distributions for POIs features within the same CBG highlight considerable variability among POIs. Figures a-c demonstrate the distributions of stars, price range, and racial segregation, respectively.
@@ Appendix B Reflective Coding Algorithm
Figure 6: Case study.
@@
1:Input: A set $O$ of tuple (POI II, POI name NN, POI reviews and images $M$, real visitor composition $\tau_{real}$, local residential composition $l_{res}$)
2:Output: Codebook $B$
3:Initialize an empty list $L$
4:for each tuple (II, NN, $M$, $tau_{real}$, $l_{res}$) in $O$ do
5:  $L_{1}\leftarrow\text{analyze}(N)$
6:  $L_{2}\leftarrow\text{analyze}(M)$
7:  $L_{\text{sum}}\leftarrow\text{summary}(L_{1},L_{2})$
8:  $\tau_{\text{diff}}\leftarrow\text{compare}(\tau_{real},l_{res})$
9:  $L_{I}\leftarrow\text{refinement}(\tau_{\text{diff}},L_{\text{sum}})$
10:  Add $L_{I}$ to $L$
11:end for
12:Initialize an empty list $B$
13:Divide $L$ into multiple subsets $L_{\text{sub}}$
14:for each $L_{\text{sub}}$ in $L$ do
15:  Initialize an empty list $B_{\text{sub}}$
16:  for each $L_{I}$ in $L_{\text{sub}}$ do
17:   $(c,e)\leftarrow\text{extract\_code\_and\_explanations}(L_{I})$
18:   Add $(c,e)$ to $B_{\text{sub}}$
19:  end for
20:  $B_{\text{sub}}\leftarrow\text{semantic\_combination}(B_{\text{sub}})$
21:  $B_{\text{sub}}\leftarrow\text{quality\_based\_removal}(B_{\text{sub}})$
22:  $B_{\text{sub}}\leftarrow\text{explanation\_refinement}(B_{\text{sub}})$
23:  $B\leftarrow\text{update}(B,B_{\text{sub}})$
24:end for
25:Return $B$
Appendix C Implementation Details
To obtain a representative data subset for the coding process, we categorize all POIs into 36 types according to the different combinations of visitor-residential racial composition gaps (e.g., higher black and asian ratios with lower white, hispanic, and other ratios). We then perform stratified sampling to randomly select POIs from each type to ensure representiveness.
For the reasoning channel, we train an MLP with hidden dimensions of 512, 128, and 64, removing the output layer to form the Rating Adapter. For the embedding channel, we fine-tune the last two layers of the pre-trained GTE-base model and extend it with three fully connected layers (512, 256, 128) to form the Embedding Adapter. For the population channel, we train an MLP with dimensions of 100, 30, and 10, extracting the weights of the first two layers to form the Population Adapter. For the neighbor-aware multi-view predictor, we adopt Graph Attention Networks (GATs) Velickovic et al. [2017] as the neighbor aggregators. The Reasoning GAT maps feature from 64 to 128, the Embedding GAT maintains a 128 to 128 mapping, and the Population GAT maintains a 30 to 30 mapping. The multi-view fusion module comprises three fully connected layers with dimensions 512, 128, and 64. During training, we first optimize the three Adapters, then freeze them before training the predictor.
During the development of REEM, we conduct a compact grid search over the principal hyper-parameters using the Philadelphia dataset. Specifically, learning rates for both the GAT and fully-connected components are explored in the interval $5\times 10^{-6},\ 5\times 10^{-3}]$, weight-decay coefficients are varied from $1\times 10^{-5}$ to $5\times 10^{-4}$, and the early-stopping patience is examined between $10$ and $25$ epochs. The width of every MLP hidden layer is tuned by doubling or halving around the final choices (approximately 6464512512 units). The configuration achieving the lowest validation MSE is selected and is released as the default value in our open-source code.
All experiments are conducted on a Linux server with the following hardware and software specifications:
  • CPU: Intel(R) Xeon(R) Platinum 8358 CPU @ 2.60GHz
[…]
  • Memory: 503.5 GB
  • Operating System: Ubuntu 22.04.3 LTS
  • Software:
    • Python 3.11.5
    • PyTorch 2.4.0 (with CUDA 12.1 support)
    • Other libraries: NumPy 1.26.0, vllm 0.6.1.post2
Appendix D Case study
@@
Appendix E Ethics Statement
Appendix F Code and Data Availability
All source code used in this work, including LLM prompts, hyperparameter configurations, data preprocessing scripts, and training scripts, is available [^8] https://github.com/LinChen-65/REEM. Given that the datasets are publicly accessible and of substantial size, we provide detailed instructions for accessing the required data, rather than including the raw data directly.

### v4

@@ 6. User Study
Figure 3. Human prediction accuracy with different information availability. Accuracy remains near random (50%) when only POI metadata (Baseline Estimation) or raw reviews (Review-informed Prediction) are available, but increases substantially once participants receive codebook-guided LLM summaries (Summary-enhanced Prediction), demonstrating the cognitive benefit of structured review distillation.
Figure 4. User preference between codebook-guided summary and vanilla summary. Across all sampled POIs, a strong majority of participants favor codebook-guided summaries, indicating that structured summaries better support human understanding of POI inclusiveness.
@@
Figure 4 presents the results of the summary preference evaluation. Across all sampled POIs, 60%-80% of participants prefer the codebook-guided summary over the vanilla one, with an average improvement of 40%, showing that our approach more effectively distills useful information to support human judgment. As Appendix  E illustrates, the codebook-guided summaries capture nuanced insights on the potential deterrants for certain minority groups, which are often absent or less explicit without codebook guidance.
7. Prediction Experiments
7.1. Experiment Settings
We compare our model against five baselines. The first baseline is an MLP making use of only the population information. The other four baselines process the social media content with a frozen pre-trained text embedding model, and fuse it with population information. We adopt four widely-used powerful embedding models: GTE-base (Li et al., 2023), BERT-base (Devlin et al., 2019), GloVE-330B (Pennington et al., 2014) and Qwen3-Embedding-8B (Zhang et al., 2025). The structures of the population and embedding MLPs are identical to those in our model. We use four metrics to evaluate model performances: mean squared error (MSE), root mean squared error (RMSE), mean absolute error (MAE), and coefficient of determination (R2).
We employ GPT-4o for reflective coding on a small subset (n=190) of POIs, and GPT-4o-mini for codebook-guided reasoning on all POIs for economic considerations. All models are trained using the Adam optimizer to minimize the MSE loss, with hyperparameters (e.g., learning rate, weight decay) tuned based on validation performances. POIs in each city are randomly split into training, validation, and test sets with a ratio of 6:2:2. For further implementation details, please refer to Appendix C and our released code. All LLM prompts used in this work are also provided in the repository.
Table 3. Ablation study of RE’EM components. Using a single channel (row 1-3) or removing the codebook (row 4) leads to performance degradation. The full model consistently achieves the best results, highlighting the complementary contributions of reasoning, embedding, and population features.
@@ 7.1. Experiment Settings
Table 4. Cross-city generalization performances using codebook derived from Philadelphia. RE’EM outperforms all baselines across Tucson, Tampa, and New Orleans, demonstrating strong transferability.
@@
Qwen3-Embedding-8B 0.0103±0.0005 0.1013±0.0028 0.0827±0.0029 0.1771±0.0106
RE’EM 0.0087±0.0005 0.0929±0.0028 0.0745±0.0024 0.2781±0.0424
7.2. Experiment Results
Table 2 presents the performances of our RE’EM model alongside four baselines, with results reported as the mean and standard deviation over five repetitions on 50% randomly sampled test sets. A key insight is that all models incorporating social media data outperform the population-only baseline. This result underscores the predictive power of social media content, which provides crucial insights beyond static population distributions. It supports our hypothesis that segregation patterns are shaped not only by demographic distributions but also by the social and cultural perceptions reinforcing the “invisible walls” in cities. Further, RE’EM consistently and significantly surpasses all embedding-based baselines. It achieves a 9.33%, 4.75%, and 3.07% improvement in MSE, RMSE, and MAE, respectively, over the population-only baseline. R2 increases to 0.3885, marking a substantial 22.79% improvement. These results indicate that while social media data encapsulates rich and valuable information, a naive embedding strategy is insufficient to fully capture the interplay between population distribution, POI attributes, and visitor dynamics. In contrast, the structured multi-view approach of RE’EM effectively integrates heterogeneous information sources, enabling accurate predictions of segregation. We provide further analyses in Appendix  D, showing that RE’EM remains stable across different base LLMs, and direct LLM prompting performs substantially worse than RE‘EM architecture.
To validate our design of different model components, we perform an ablation study. As Table 3 shows, each channel (reasoning, embedding, population) exhibits certain predictive capabilities, but none achieves optimal performance. Without the codebook, relying solely on vanilla LLM rating yields markedly worse performance. The full RE’EM model, which strategically combines all three channels and leverages the codebook for structured reasoning, consistently delivers superior results. Thus, the integration of multi-view signals enables complementary information flow, allowing the model to refine its predictions beyond what any single feature source can achieve.
To assess the generalization capability across cities, we use the codebook obtained in Philadelphia to train models in the other three cities. As shown in Table 4, RE’EM consistently achieves the best performance in all three cities along all metrics. Compared to the strongest baseline, RE’EM achieves substantial improvements, with MSE reduced by as much as 12.12%, MAE reduced by as much as 9.17% and R² increasing by as much as 35.99%. These findings indicate that RE’EM, and specifically, our constructed codebook, captures fundamental segregation mechanisms that extend beyond city-specific characteristics, making it a promising tool across diverse geographic and socio-economic settings.
8. Discussion
@@ 9. Conclusion
10. Ethics Statement
Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China (No.62472241) and the National Key Research and Development Program of China (2024YFC3307605).
@@ Appendix A Variability of POI features
Figure 5. The Coefficient of Variation distributions for POIs features within the same CBG highlight considerable variability among POIs. Figures a-c demonstrate the distributions of stars, price range, and racial segregation, respectively.
@@ Appendix B Reflective Coding Algorithm
Appendix C Implementation Details
To obtain a representative subset for the coding process, we categorize all POIs into 36 types base on different combinations of visitor-residential composition gaps (e.g., higher black and asian ratios with lower white, hispanic, and other ratios) and perform stratified sampling within each type to ensure diversity.
For the reasoning channel, we train an MLP with hidden dimensions of 512, 128, and 64, removing the output layer to form the Rating Adapter. For the embedding channel, we fine-tune the last two layers of the pre-trained GTE-base model and extend it with three fully connected layers (512, 256, 128) to form the Embedding Adapter. For the population channel, we train an MLP with dimensions of 100, 30, and 10, extracting the weights of the first two layers to form the Population Adapter. For the neighbor-aware multi-view predictor, we adopt Graph Attention Networks (GATs) (Velickovic et al., 2017) as the neighbor aggregators. The Reasoning GAT maps feature from 64 to 128, the Embedding GAT maintains a 128 to 128 mapping, and the Population GAT maintains a 30 to 30 mapping. The multi-view fusion module comprises three fully connected layers with dimensions 512, 128, and 64. During training, we first optimize the three Adapters, then freeze them before training the predictor.
We conduct a compact grid search on the Philadelphia dataset, exploring learning rates in $5\times 10^{-6},5\times 10^{-3}]$, weight decays in $10^{-5},5\times 10^{-4}]$, early-stopping patience between 10 and 25. The width of every MLP hidden layer is tuned by doubling or halving around the final choices (approximately 6464512512 units). The configuration achieving the lowest validation MSE is selected and is released as the default value in our open-source code.
[…]
@@ Appendix C Implementation Details
1:Set of POIs $\mathcal{I}$ with for each POI $i\in\mathcal{I}$:
2: - Social media content $c_{i}=(\mathcal{N}_{i},\{\text{review}_{k},\text{image}_{k}\}_{k=1}^{m_{i}})$
3: - Observed visitor composition $\boldsymbol{\tau}_{i}=[\tau_{iq}]_{q=1}^{Q}$ (Q racial groups)
4: - Local residential composition $\mathbf{p}_{i}=[p_{iq}]_{q=1}^{Q}$
5:Codebook $\mathcal{B}=\{(d_{j},e_{j})\}_{j=1}^{M}$ with dimensions $d_{j}$ and explanations $e_{j}$
6:// Phase 1: Reflective Attributor
7:Initialize insight set $\mathcal{L}\leftarrow\emptyset$
8:for each POI $i\in\mathcal{I}$ do
9:  // Multi-source insight generation
10:  $\mathcal{L}_{i}^{\text{name}}\leftarrow\text{LLM}_{\text{analyze}}(\mathcal{N}_{i})$ $\triangleright$ Extract name signals
11:  $\mathcal{L}_{i}^{\text{text}}\leftarrow\text{LLM}_{\text{analyze}}(\{\text{review}_{k}\})$ $\triangleright$ Analyze textual reviews
12:  $\mathcal{L}_{i}^{\text{img}}\leftarrow\text{LLM}_{\text{analyze}}(\{\text{image}_{k}\})$ $\triangleright$ Analyze visual content
13:  $\mathcal{L}_{i}^{\text{sum}}\leftarrow\text{LLM}_{\text{synthesize}}(\mathcal{L}_{i}^{\text{name}},\mathcal{L}_{i}^{\text{text}},\mathcal{L}_{i}^{\text{img}})$
14:  // Reflection-guided refinement via abductive reasoning
15:  $\Delta\boldsymbol{\tau}_{i}\leftarrow\boldsymbol{\tau}_{i}-\mathbf{p}_{i}$ $\triangleright$ Compute demographic discrepancy
16:  $\mathcal{L}_{i}^{\text{refined}}\leftarrow\text{LLM}_{\text{reflect}}(\mathcal{L}_{i}^{\text{sum}},\Delta\boldsymbol{\tau}_{i})$
17:  $\mathcal{L}[i]\leftarrow\mathcal{L}_{i}^{\text{refined}}$
18:end for
19:// Phase 2: Code Summarizer
20:Partition $\mathcal{L}$ into $K$ stratified subsets: $\{\mathcal{L}_{\text{sub}}^{(k)}\}_{k=1}^{K}$
21:Initialize codebook $\mathcal{B}\leftarrow\emptyset$
22:for each subset $\mathcal{L}_{\text{sub}}$ do
23:  // Extract candidate codes
24:  $\mathcal{C}_{\text{sub}}\leftarrow\emptyset$
25:  for each insight $\ell\in\mathcal{L}_{\text{sub}}$ do
26:   $(c,e)\leftarrow\text{LLM}_{\text{extract}}(\ell)$
27:   $\mathcal{C}_{\text{sub}}\leftarrow\mathcal{C}_{\text{sub}}\cup\{(c,e)\}$
28:  end for
29:  // Iterative refinement on combined set
30:  $\mathcal{B}\leftarrow\mathcal{B}\cup\mathcal{C}_{\text{sub}}$ $\triangleright$ Merge candidates with existing codebook
31:  $\mathcal{B}\leftarrow\text{LLM}_{\text{semantic}}(\mathcal{B})$ $\triangleright$ LLM merges redundant dimensions
32:  $\mathcal{B}\leftarrow\text{LLM}_{\text{quality}}(\mathcal{B})$ $\triangleright$ LLM filters low-generalizability codes
33:  $\mathcal{B}\leftarrow\text{LLM}_{\text{refine}}(\mathcal{B})$ $\triangleright$ LLM generates thematic descriptions
34:end for
35:return $\mathcal{B}$
Table 5. Evaluation of different base LLMs in the RE’EM framework. Replacing GPT-4o-mini with other SOTA LLMs yields nearly identical results, showing that RE’EM is robust to the choice of base model.
Base LLM MSE↓ RMSE↓ MAE↓ R2
GPT-4o-mini 0.0068±0.0003 0.0823±0.0021 0.0662±0.0021 0.3885±0.0278
Qwen-plus 0.0068±0.0003 0.0822±0.0021 0.0661±0.0021 0.3905±0.0289
Llama4-scout 0.0068±0.0004 0.0823±0.0022 0.0660±0.0021 0.3896±0.0305
DeepSeek-v3.2-exp 0.0068±0.0004 0.0823±0.0022 0.0662±0.0021 0.3883±0.0301
Table 6. Comparison with direct LLM prompting baseline. Removing RE’EM pipeline leads to severe performance degradation, confirming the necessity of RE’EM.
Model MSE↓ RMSE↓ MAE↓ R2
Qwen-plus 0.1281±0.0039 0.3579±0.0054 0.3202±0.0068 -10.7420±0.4472
Llama4-scout 0.0586±0.0020 0.2420±0.0043 0.1916±0.0042 -4.3700±0.2091
DeepSeek-v3.2-exp 0.0689±0.0022 0.2625±0.0042 0.2127±0.0030 -5.3191±0.2724
RE’EM 0.0068±0.0003 0.0823±0.0021 0.0662±0.0021 0.3885±0.0278
Figure 6. Case study comparing the vanilla summary and codebook-guided summary for the same venue. The codebook-guided summary captures more nuanced factors influencing demographic appeal, which are absent in the vanilla summary, illustrating the interpretability and analytical advantage of structured reasoning.
Appendix D Additional Evaluation on Different LLMs
To further examine whether RE’EM depends on a particular language model, we conduct two additional evaluations involving both open- and closed-source state-of-the-art LLMs. First, we replace the base model (GPT-4o-mini) used in the reasoning channel with three strong alternatives: Qwen-plus (Yang et al., 2025), Llama4-scout (AI, 2025), and DeepSeek-v3.2-exp (DeepSeek-AI, 2025), keeping all other components unchanged. As shown in Table 5, the performance across all metrics remains nearly identical, indicating that RE’EM is highly robust to the choice of the base LLM.
Second, we assess whether the performance gain comes from the RE’EM architecture rather than the inherent ability of LLMs. We directly prompt the same LLMs to predict segregation from social media content and population distribution, removing the RE’EM pipeline. Table 6 shows a dramatic drop in performance for all models, confirming that naive LLM prompting is insufficient and that the structured multi-channel design of RE’EM is essential for reliable quantitative prediction.
Appendix E Case study